# GenTorrent: Scaling Large Language Model Serving with An Overlay Network

### v3

@@ 7 Conclusion
This work introduces GenTorrent, a novel decentralized infrastructure designed for scalable and privacy-preserving LLM serving. We present designs addressing four key challenges: overlay network organization, anonymous communication, overlay forwarding for efficiency, and verification of model serving quality. These issues are fundamental and inherent to any future decentralized LLM-serving system. We implement a prototype of GenTorrent and evaluation results show that GenTorrent achieves low serving latency with minimal overhead for security and trustworthy features. In the future, we plan to deploy GenTorrent at scale and encourage broader community engagement in the research and use of GenTorrent. We believe this work offers a new direction that will attract further research and eventually help democratize future AI innovations and deployment.
@@
Limitations of overlay Onion routing. Onion routing [36] was proposed to achieve anonymous routing in the network layer where the forwarding nodes are network routers, such as Tor [13]. If we apply the same idea to the overlay network in GenTorrent, the main challenge is that the failure probability of an Onion path increases exponentially with the path length, especially for an overlay network of users with a high churn rate. In addition, each relay needs to perform public-key descryption on the entire message, introducing non-trivial overhead for users.
Strengths and weaknesses of slicing protocols. One idea is to use sliced messages for overlay anonymity [22, 34]. In garlic routing [34], the sender sends $n$ slices, each is about $1/k$ of the original message size long and is called a clove. If the receiver gets $k$ out of $n$ cloves, it can recover the original message. Garlic achieves 1) anonymity without encryption and decryption operations on the paths; 2) reliability under node failures; 3) the bandwidth cost does not increase significantly with the nkn-k redundant cloves. Garlic routing has a few weaknesses if we apply it directly to GenTorrent. 1) Garlic relies on random walks to establish routing paths and the corresponding “proxies”. However, random walks add uncertainty to the success of finding enough proxies. 2) Garlic preserves the provider (model node) anonymity, which is unnecessary for GenTorrent.
@@
Determining the optimal configuration for the chunk length array $L$ presents a significant challenge. Setting the chunk sizes too small results in a HashRadix Tree (HR-Tree) with excessive depth, potentially over-differentiating prefixes. Conversely, if chunk sizes are too large, requests with substantial prefix overlap and high cache affinity might be incorrectly treated as entirely distinct. To address this, GenTorrent incorporates a Sentry module. This component is responsible for collecting incoming requests and detecting common patterns, such as system prompts, which can inform the dynamic adjustment or initial configuration of the chunking strategy defined by $L$. How to set $L$ to effciently utlize the HR-Tree? As described in Fig. 19, we define $S=s_{1},s_{2},\dots,s_{n}$ be the lengths of the distinct common system prompts detected by the Sentries, sorted in ascending order. The Sentry module uses these lengths, along with a small, fixed separator length $\delta$, to construct the chunk length array L used in Algorithm 1. The array L is defined as follows:
@@
A4. Definition of Anonymity Metric

### v4

@@ 8 Conclusion
This work introduces PlanetServe, a novel decentralized infrastructure designed for scalable and privacy-preserving LLM serving. We present designs addressing four key challenges: overlay network organization, anonymous communication, overlay forwarding for efficiency, and verification of model serving quality. These issues are fundamental and inherent to any future decentralized LLM-serving system. We implement a prototype of PlanetServe and evaluation results show that PlanetServe achieves low serving latency with minimal overhead for security and trustworthy features. In the future, we plan to deploy PlanetServe at scale and encourage broader community engagement in the research and use of PlanetServe. We believe this work offers a new direction that will attract further research and eventually help democratize future AI innovations and deployment.
@@
Limitations of overlay Onion routing. Onion routing [Onion1998] was proposed to achieve anonymous routing in overlay networks, such as Tor [Tor04]. If we apply the same idea to the overlay network in PlanetServe, there are a few main challenges. First, the first layer of Onion nodes (called as guards in Tor) will definitely know the sending user of a request, which should be avoided in PlanetServe. Second, the failure probability of an Onion path increases exponentially with the path length, especially for an overlay network of users with a high churn rate. In addition, each relay needs to perform public-key decryption on the entire message, introducing non-trivial overhead for users.
Strengths and weaknesses of slicing protocols. One idea is to use sliced messages for overlay anonymity [IS-NSDI07, garlic]. In garlic routing [garlic], the sender sends $n$ slices, each is about $1/k$ of the original message size long and is called a clove. If the receiver gets $k$ out of $n$ cloves, it can recover the original message. Garlic achieves 1) anonymity without encryption and decryption operations on the paths; 2) reliability under node failures; 3) the bandwidth cost does not increase significantly with the nkn-k redundant cloves. Garlic routing has a few weaknesses if we apply it directly to PlanetServe. 1) Garlic relies on random walks to establish routing paths and the corresponding “proxies”. However, random walks add uncertainty to the success of finding enough proxies. 2) Garlic preserves the provider (model node) anonymity, which is unnecessary for PlanetServe.
@@
Determining the optimal configuration for the chunk length array $L$ presents a significant challenge. Setting the chunk sizes too small results in a HashRadix Tree (HR-Tree) with excessive depth, potentially over-differentiating prefixes. Conversely, if chunk sizes are too large, requests with substantial prefix overlap and high cache affinity might be incorrectly treated as entirely distinct. To address this, PlanetServe incorporates a Sentry module. This component is responsible for collecting incoming requests and detecting common patterns, such as system prompts, which can inform the dynamic adjustment or initial configuration of the chunking strategy defined by $L$. How to set $L$ to effciently utlize the HR-Tree? As described in Fig. 18, we define $S=s_{1},s_{2},\dots,s_{n}$ be the lengths of the distinct common system prompts detected by the Sentries, sorted in ascending order. The Sentry module uses these lengths, along with a small, fixed separator length $\delta$, to construct the chunk length array L used in Algorithm 1. The array L is defined as follows:
@@
A4. Analysis of choice of $n$ and $k$ in anonymous routing.
In anonymous routing of PlanetServe, we need $k$ of $n$ cloves to arrive for a successful communication. There are $l=3$ relays on each path. Assume the failure rate of an overlay node during the time of a round of communication is $f$. The probably that one path succeeds is $(1-f)^{3}$. The rate that at least $k$ out $n$ succeed is:
$P(X\geq k)=\sum_{i=k}^{n}\binom{n}{i}(1-f)^{3i}(1-(1-f)^{3})^{n-i}$
Using $n=4$ and $k=3$, even with a failure rate as high as $3\%$, the success rate is $95\%$.
A5. Definition of Anonymity Metric
@@
Figure 19: Computation cost for HR-tree update
Figure 20: Network cost for HR-tree update
Figure 21: Latency of LLM serving with A1000 GPUs
A6. HR-tree Update Cost
We evaluate the computation cost and network cost of HR-tree update by comparing the full broadcast design and the proposed delta update and show the results in Fig. 19 and Fig. 20, respectively. The results show that delta update significantly reduces the cost.
A7. Latency of LLM serving with A1000 GPUs
We evaluate the latency of LLM serving with A1000 GPUs and show the results in Fig. 21. Compared to the A100 results in Fig. 14, PlanetServe shows similar advantages.
Figure 22: mixed with upper-bound
A8. Latency Comparison with Centralized Baselines
We compare PlanetServe with a centralized baseline of 8 A100 GPUs without KV-cache sharing and centralized serving with 8 A100 GPUs, running vLLM with the default continuous-batching scheduler and tensor parallelism. Fig. 22 shows the latency comparison. We find that PlanetServe’s latency is close to the Centralized sharing method, but significantly shorter than that of Centralized non-sharing.
A9. Resiliency to Intersection Attacks
An intersection attack [Intersection-CCS13] can be launched to anonymity systems if the target uses pseudonymity and the attacker has access to the user’s history of visits and ISP logs. PlanetServe treats each sequence of prompts as independent. Hence users do not use pseudonymity and the attacker cannot obtain user’s history. Even though the attacker can access ISP logs (very big assumption!), they cannot match the user behavior among the massive amount of Internet users.
A10. Real-World Routing Latency Measurements
Figure 23: Measured Session-Establish and In-Session Latency Across Regions
To evaluate the practical impact of network overhead, we measured routing latency on real cloud infrastructure. Specifically, we deployed 14 AWS t3.micro instances, with each hop placed in a different region (datacenter), and used a mixed workload to simulate real prompt payloads. Over the course of 4000 runs, we observed three failures in the across-USA setting and one failure in the across-world setting.
Results are summarized in Fig. 23. In the across-USA experiment (four regions in the United States), the session-establishment latency was 168.9 ms (P99: 256.8 ms), while the steady in-session latency was 92.9 ms (P99: 179.2 ms). In the across-world experiment (five regions across North America, Asia, Europe, and South America), the session-establishment latency was 577.4 ms (P99: 685.8 ms), and the steady in-session latency was 919.6 ms (P99: 1025.5 ms).
Even at inter-continental scales, these latencies are modest compared to LLM inference time (see Fig. 14), suggesting that communication does not dominate end-to-end serving time.